OpenStudy (anonymous):

Alphonse is training for a marathon. He runs 6 1/2 miles every 2/3 hour. How far does he run in 2 hours?

OpenStudy (anonymous):

You are given a distance, and you are given a time. So you know how many miles he is running in a certain amount of time

OpenStudy (anonymous):

how do i show how to solve this question in unit rate?

OpenStudy (wolf1728):

Multiply 6.5 miles every 2/3 hour by 3 will give miles per 2 hours. 6.5 * 3 = 19.5 miles / 2 hours.

OpenStudy (wolf1728):

mody - by unit rate do you mean converting the velocity to miles per hour?

OpenStudy (anonymous):

I don't think so. we are converting fractions to unit rates.

OpenStudy (wolf1728):

and the unit rates would be?

OpenStudy (anonymous):

that's what Im trying to find out by the original question asked.

OpenStudy (wolf1728):

So, could you post the original question?

OpenStudy (anonymous):

Alphonse is training for a marathon. He runs 6 1/2 miles every 2/3 hour. How far does he run in 2 hours?

OpenStudy (wolf1728):

So , that is the original question. Anyway, where are you asked about converting everything to unit rates?

OpenStudy (anonymous):

we are doing unit rate/fractions. Alphonse is training for a marathon. He runs 6 1/2 miles every 2/3 hour. How far does he run in 2 hours? we need to show our work on how you multiply and divide fractions to get the unit rate answer.

OpenStudy (wolf1728):

Well if we multiply 6 1/2 by 3/2, it will give you miles per hour.

OpenStudy (anonymous):

how do we do it by dividing?

OpenStudy (wolf1728):

Does it really require dividing? If you are supposed to learn about fractions, I'm guessing you have other problems to solve? If that's the case, maybe fraction division is in one of those problems.

OpenStudy (wolf1728):

Well anyway, converting 6 1/2 miles per 2/3 hour: Let's convert 6 1/2 miles to halves and that's 13/2 Next we multiply this by 3/2 and we have 13/2 * 3/2 = 39/4 = 9 (3/4) miles per hour.

OpenStudy (wolf1728):

For 2 hours, we multiply by 2: 39/4 * 2 = 78/4 = 19 2/4 miles = 19 1/2 miles

OpenStudy (wolf1728):

Hope that helps you out.

OpenStudy (anonymous):

how do i award a medal?

OpenStudy (anonymous):

You were a great help Wolf. Thank You!!!